what must be the the minimum initial velocity to reach 1000m range? what is the total time required to reach this range


1
Expert's answer
2021-10-24T18:26:19-0400

Given:

"R=1000\:\rm m"


The range

"R=\frac{v_0^2\sin2\theta}{g}"

The minimum speed corresponds to the angle of throws "\theta=45^{\circ}" and equal

"v_0=\sqrt{gR}=\sqrt{9.8*1000}=99\:\rm m/s"